### v1

# Homomorphism counting for immersion-closed classes is not isomorphism

Abstract
Lovász proved that two graphs $G$ and $H$ are isomorphic if $\hom(K,G)=\hom(K,H)$ for all graphs $K$, where $\hom(G_{1},G_{2})$ denotes the number of homomorphisms from $G_{1}$ to $G_{2}$. Dvořák showed that it suffices to count homomorphisms from all $2$-degenerate graphs $K$. On the other hand, for several interesting graph classes $\mathcal{M}$, it has been shown that there exist non-isomorphic graphs $G$ and $H$ such that $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}$. Most such classes are minor-closed classes and Roberson conjectured that every proper minor-closed and union-closed graph class $\mathcal{M}$ has the property of there existing non-isomorphic graphs that are indistinguishable by homomorphism counts from $\mathcal{M}$. There has been an effort to prove Roberson’s conjecture as it is believed that minor-closed classes play a special role in the context of homomorphism indistinguishability.
We show that this special role, if so, must be shared, by proving an analogue of Roberson’s conjecture holds for a rich family of non-minor-closed classes. Namely, we prove that for any proper immersion-closed and union-closed class $\mathcal{M}$, there exist non-isomorphic graphs $G$ and $H$ such that $\hom(K,G)=\hom(K,H)$ for all $K\in\mathcal{M}$. This extends a result of Roberson on homomorphism indistinguishability over bounded degree graphs, and gives an almost full picture since our result cannot be extended in the natural way, that is, by replacing immersions with topological minors, due to a result of Neuen and Seppelt.
Our main result is obtained as a consequence of a colouring result which may be of independent interest: Every graph with a ${t\choose 2}(7t+7)$-oddomorphism admits a $K_{t}$-immersion. We also give an alternative proof of the known result that homomorphism counts from graphs of bounded treewidth does not distinguish non-isomorphic graphs. This again follows from a new result on oddomorphisms: Every graph with a $(t+2)$-oddomorphism has treewidth at least $t+1$. This latter bound is best possible.
1 Introduction
Determining the complexity of graph isomorphism is one of the most important open problems in theoretical computer science. Currently, the best known worst-case running time of a graph isomorphism algorithm is quasi-polynomial due to Babai [3]. Babai’s algorithm uses a combination of sophisticated algebraic techniques as well as the $k$-dimensional Weisfeiler-Leman algorithm for small values of $k$. Practical algorithms, such as NAUTY, all use a framework called individualization-refinement [18]. The solvers all build on a framework called individualization-refinement. While we will not need the precise details, they all revolve around finding fast “refinement” algorithms, and then getting them “unstuck” if needed with individualization arguments.
The most well known of these refinement algorithm is called colour refinement or the one-dimensional Weisfeiler-Leman algorithm. This is a practical graph isomorphism solver and it was shown in [1] that, in a smoothed analysis setting, colour refinement solves the majority of cases of graph isomorphism in near-linear time. While the algorithmic definitions of colour refinement are more useful for actual computation, in this note we will instead consider the homomorphism definition.
Recall that for two graphs $G$, $H$, a graph homomorphism from $G$ to $H$ is a function $f:V(G)\to V(H)$ such that for all $e=xy\in E(G)$, we have $f(x)f(y)\in E(H)$. Let $\hom(G,H)$ denote the number of homomorphisms from $G$ to $H$.
For two graphs $G$ and $H$, we say that the colour refinement algorithm fails to distinguish $G$ from $H$ if for all trees $T$, we have $\hom(T,G)=\hom(T,H)$. For those used to the algorithmic version of colour refinement, this is perhaps a surprising equivalence, which was shown by Dvořák in [9]. Nevertheless, we take this as a definition of the colour refinement.
While colour refinement is extremely effective in practice, it also fails on relatively simple classes of graphs. For example, it is easy to see it cannot distinguish two $d$-regular graphs on the same number of vertices. In particular, it fails to distinguish $2K_{3}$ from $C_{6}$, i.e., for every tree $T$, we have $\hom(T,2K_{3})=\hom(T,C_{6})$. Therefore, a natural approach to extending the colour refinement algorithm is to count homomorphisms from wider classes of graphs beyond trees. This produces spectacular results and often maintains the efficiency of colour refinement.
Given a class $\mathcal{M}$ of graphs, we say that two graphs $G$ and $H$ are $\mathcal{M}$-isomorphic, denoted $G=_{\mathcal{M}}H$, if $\hom(K,G)=\hom(K,H)$ for all graphs $K\in\mathcal{M}$. Lovász famously proved:
Theorem 1.1 ([16]).
If $\mathcal{M}$ is the class of all graphs, then $G=_{\mathcal{M}}H$ if and only if $G$ and $H$ are isomorphic.
Thus, homomorphism counts from trees produce the colour refinement algorithm and homomorphism counts from all graphs produce graph isomorphism. Homomorphism counts from other classes in between produce interesting results as well. Table 1 shows some of the various relationships, but is by no means exhaustive.
Graph class $\mathcal{M}$ $G=_{\mathcal{M}}H$ Complexity
Cycles $G$ and $H$ are cospectral (see [13]) $O(n^{3})$ (see [13])
Graphs with treewidth $\leq k$ $G$ and $H$ are indistinguishable by the $k$-dimensional Weisfeiler-Leman Algorithm (see [13]) $O(n^{k+1}\log(n))$ [14])
Planar graphs $G$ and $H$ are quantum isomorphic [17] Undecidable [2]
2-degenerate graphs $G$ and $H$ are isomorphic [9] At worst quasi-polynomial time [3]
All graphs $G$ and $H$ are isomorphic [16] At worst quasi-polynomial time [3]
Treedepth $\leq k$ $G$ and $H$ admit a coKleisli isomorphism with respect to Ehrenfeucht-Fraïssé comonad [12] LOGSPACE [6]
Disjoint union of paths $G$ and $H$’s adjacency matrices have the same multisets of main eigenvalues [25] Exact Logspace [25]
$K_{5}$-topological-minor-free $G$ and $H$ are isomorphic [21] At worst quasi-polynomial time [3]
Table 1: Homomorphism counting equivalence for different graph classes. For each class $\mathcal{M}$, the table shows the characterization of $G=_{\mathcal{M}}H$ and computational complexity of the equivalence test.
We see from Table 1 that for a given graph class $\mathcal{M}$, it is not obvious what the relation $_{\mathcal{M}}$ is, nor what the complexity of determining if $G=_{\mathcal{M}}H$ is. However, important graph isomorphism tests appear by considering various natural classes $\mathcal{M}$. In Table 1 all the classes $\mathcal{M}$ for which $_{\mathcal{M}}$ differs from isomorphism are all minor-closed [^1] Given a graph $G$, a graph $H$ is a minor of $G$ if it can be obtained by deleting edges, contracting edges, and deleting isolated vertices. classes, or at least closed under edge contractions. This motivates the following conjecture of Roberson.
Conjecture 1.2 ([22]).
If $\mathcal{M}$ is a minor-closed and union-closed class that does not contain all graphs, then there are non-isomorphic graphs $G$ and $H$ such that $G=_{\mathcal{M}}H$.
However, even if this conjecture were true, it would not give the full picture. Indeed if $\mathcal{M}$ is the class of graphs with maximum degree at most $d$, then $_{\mathcal{M}}$ is not the isomorphism relation as shown in [22], and the class of graphs with maximum degree at most $d$ is not minor-closed for any fixed $d\geq 3$. So what is the correct level of generality? A natural starting point to unify bounded degree graphs and minor-closed classes are classes that are closed under topological minors [^2] GG contains $H$ as a topological minor if a subgraph of $G$ contains a subdivision of $H$. Yet recently Neuen and Seppelt [21] showed that homomorphism counting is isomorphism even for the class of graphs with no $K_{5}$ topological minor. The authors of [21] interpret this last result as evidence of the “special role” that minor-closed classes play in demarcating the graph classes $\mathcal{M}$ for which $_{\mathcal{M}}$ differs from isomorphism. While this special role can only be confirmed by the proof of Conjecture 1.2, we show that the role of immersions is at least as important by proving the following, which is the immersion-analogue of Roberson’s conjecture.
Theorem 1.3.
Let $\mathcal{M}$ be a proper immersion-closed and union-closed class. Then there are non-isomorphic graphs $G,H$ such that for all $K\in\mathcal{M}$, we have $\hom(K,G)=\hom(K,H)$.
Immersion is a well-studied [^3] See e.g. [4, 5, 8, 15, 24], as a small sampling. graph order incomparable to the minor order, which nevertheless shares important properties of it: it is a well-quasi-order, as conjectured by Nash-Williams in 1963 [19] and proved by Robertson and Seymour [23]. Moreover, if a graph $G$ contains $H$ as a topological minor, then $G$ contains $H$ both as a minor and as an immersion. In particular, this tells us that the result of Neuen and Seppelt helps to emphasize the importance of Theorem 1.3, since its natural generalisation would be to consider classes closed under topological minors. We now give the definition of immersion.
Definition 1.4.
Let $G$ be a graph, and suppose we have three distinct vertices $u,v,w$ with edges uvuv,vwvw. To split off $uv$ and $vw$ is to delete the edges $uv$ and $vw$, and add the edge $uw$ (possibly adding a parallel edge if necessary). A graph $G$ contains a graph $H$ as an immersion if $H$ can be obtained from a subgraph of $G$ by splitting off pairs of edges and deleting isolated vertices.
Note that for every positive integer $d$, the class of graphs with maximum degree at most $d$ is immersion-closed. Therefore, Theorem 1.3 generalises the result of Roberson [22] that homomorphisms from bounded degree classes does not capture isomorphism. The following further strengthening may be possible and we leave it as an open problem.
Conjecture 1.5.
Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ be two distinct immersion-closed and union-closed classes. Then for some $i,j\in\{1,2\}$ with $i\neq j$, there are graphs $G$ and $H$ such that $G=_{\mathcal{M}_{i}}H$ and $G\neq_{\mathcal{M}_{j}}H$.
To prove Theorem 1.3, we utilize tools developed by Roberson in [22], that essentially reduce the problem to a colouring one. Recall that $N_{G}(v)$ denotes the neighbourhood of $v$ in a graph $G$. The following definitions are introduced in [22].
Definition 1.6.
Let $f\colon V(G)\rightarrow S$ be a proper colouring of a graph $G$. A vertex $u\in V(G)$ is $f$-odd (respectively, $f$-even) if $N_{G}(v)$ contains an odd (respectively, even) number of vertices of each colour in $S\setminus\{f(v)\}$.
Of course, it is possible for a vertex to be neither odd nor even.
Definition 1.7.
A proper $t$-colouring $f$ of a graph $G$ is a $t$-oddomorphism, if
  1. 1.
    Every vertex of $G$ is either $f$-even or $f$-odd.
  2. 2.
    Every colour class contains an odd number of odd vertices.
If $G$ admits a $t$-oddomorphism, we write $G\xrightarrow{\text{odd}}K_{t}$. In Section 3 we prove the following.
Theorem 1.8.
If $G\xrightarrow{\text{odd}}K_{{t\choose 2}(7t+7)}$, then $G$ contains a $K_{t}$-immersion.
While it does not appear at all related, it turns out that Theorem 1.8 and a result in [22] implies Theorem 1.3. We show that this is the case in Section 5. The bound we obtain is not necessarily optimal, and we make the following conjecture.
Conjecture 1.9.
There exists an absolute constant $c$ such that if $G\xrightarrow{\text{odd}}K_{ct}$, then $G$ contains a $K_{t}$-immersion.
It might even be possible that $c=1$ can be taken. In fact, Theorem 1.10 below gives that $c=1$ can indeed be taken when $t=2,3$. This theorem, moreover, is a tight lower bound on the treewidth of graphs that have a $t$-oddomorphism. (We recall the definition of treewidth in Section 2.3.)
Theorem 1.10.
If $f:G\xrightarrow{\text{odd}}K_{t+2}$, then $\operatorname{tw}(G)\geq t+1$.
The $t+1$ bound given here is best possible since $K_{t+2}\xrightarrow{\text{odd}}K_{t+2}$ and $\operatorname{tw}(K_{t+2})=t+1$. Theorem 1.10 also gives an alternative proof that homomorphism counts from graphs of bounded treewidth graph do not fully capture isomorphism. However, as we have mentioned above, homomorphism counts from bounded treewidth graphs is exactly described by the $k$-dimensional Weisfeiler-Leman algorithm, which is already known to not fully capture isomorphism [20].
The rest of the paper is structured as follows. In Section 2.3 we give some definitions, state a known key result, and prove some preliminary lemmas about oddomorphisms. In Section 3, we prove Theorem 1.8. In Section 4, we prove Theorem 1.10. Finally, in Section 5, we prove Theorem 1.3. We devote the rest of this section to sketch the proofs of our oddomorphism results.
To prove Theorem 1.8 the idea is as follows. We observe first that if $f$ is a ${t\choose 2}(7t+7)$-oddomorphism, then if there are any $2$-coloured cycles, deleting the edges of the cycle preserves having an oddomorphism. We therefore may assume that the union of any two colour classes is a forest. Next we observe if the union of two colour classes has a path from an $f$-odd vertex in one colour to an $f$-odd vertex in another colour, then we can split off the entire path and still have an ${t\choose 2}(7t+7)$-oddomorphism. Thus the structure of a minimal counterexample is that between any two colour classes, say $C_{i}$ and $C_{j}$, in a component that contains an $f$-odd vertex $v$ say in $C_{i}$, we have that either $v$ is adjacent to an $f$-odd vertex in $C_{j}$, or there is a $2$-coloured path $P$ in $G[C_{i},C_{j}]$ with endpoints $v$ and another $f$-odd vertex in $C_{i}$. In a systematic way, we split off all these paths, and get a new graph $G^{\prime}$ (which need not have an ${t\choose 2}(7t+7)$-oddomorphism).
Utilizing the $f$-odd properties of oddomorphisms, there are two possibilities for $G^{\prime}$. Either there will be two vertices given the same colour $x,y$ which have at least $t\choose 2$} parallel edges between them, or the graph obtained after removing parallel edges has minimum degree at least $7t+7$. In the latter case, a theorem of Gauthier, Le and Wollan [11] gives the $K_{t}$-immersion.
In the former case, we have two vertices with the same colour that have at least $t\choose 2$} edge disjoint $2$-coloured paths between them, and identifying these vertices, deleting the $2$-coloured paths, and removing the symmetric difference of the neighbourhoods, gives a smaller graph with an ${t\choose 2}(7t+7)$-oddomorphism, and we show that the $K_{t}$-immersion of the smaller graph lifts to an immersion of the original graph.
To prove Theorem 1.10 the idea is the following. Suppose we have a $(t+2)$-oddomorphism and our graph has treewidth at most $t$, and that we have a vertex minimal counterexample. Take an optimal tree decomposition which we can assume has some nice “smooth” properties, and root the tree decomposition at some vertex. By using a breadth first search from the root, find an odd vertex which appears last (so we count from the root each the first time each odd vertex appears in a bag, and we take the last odd vertex to appear). Since we have a $(t+2)$-oddomorphism, and each bag has size at most $t+1$, this odd vertex will have a neighbour in a bag farther away from the root, and further the colour of this neighbour will not appear in its current bag. Then by looking at a Kempe chain between the two colours of the odd vertex and this other vertex lower in the tree, we will find two vertices with the same colour that we can identify and keep the treewidth the same, and preserve the oddomorphism. In this way we find a smaller counterexample, which is a contradiction.
2 Preliminaries
2.1 An equivalent definition for immersions and a key result
An equivalent (see [10]) and particularly useful definition of immersions reads as follows. A graph $G$ contains a graph $H$ as an immersion if there is a one-to-one mapping $f:V(H)\to V(G)$ and a collection of edge-disjoint paths in $G$, one for each edge of $H$, such that the $(u,v)$-path $P_{uv}$ corresponding to the edge $uv\in E(H)$ has endpoints $f(u)$ and $f(v)$. The vertices in $f(V(H))$ are called branch vertices.
We will need the following theorem of Gauthier, Le and Wollan. (See [10] and [7] for previous results on minimum degree conditions forcing clique immersions.)
Theorem 2.1 ([11]).
Every simple graph with minimum degree $7t+7$ has a $K_{t}$-immersion.
2.2 Definition of treewidth
Let $G$ be a graph. A tree decomposition of $G$ is a pair $\left(T,\{\beta_{t}\}_{t\in V(T)}\right)$ where $T$ is a tree and each $\beta_{t}\subseteq V$ (called a bag) such that:
  1. 1.
    $\bigcup_{t\in V(T)}\beta_{t}=V$;
  2. 2.
    for every edge $uv\in E(G)$, there exists $t\in V(T)$ with $uv\in E(G[\beta_{t}])$;
  3. 3.
    for every vertex $v\in V(G)$, the set $\{t\in V(T)\mid v\in\beta_{t}\}$ induces a connected subtree of $T$.
The width of a tree decomposition $\left(T,\{\beta_{t}\}_{t\in V(T)}\right)$ is $\max_{t\in V(T)}|\beta_{t}|-1$. The treewidth of a graph $G$, denoted $\operatorname{tw}(G)$, is the minimum width over all tree decompositions of $G$.
2.3 Preliminaries about oddomorphisms
First we record a property that is immediate from the definition, and that we will use implicitly throughout.
Proposition 2.2.
Let $f:G\xrightarrow{\text{odd}}K_{n}$. Then every $f$-odd vertex has degree at least n1n-1.
Proof.
An $f$-odd vertex has at least one neighbour in each colour class other than its own colour, and hence has degree at least n1n-1. ∎
Now we give an important operation which preserves oddomorphisms.
Definition 2.3.
Given an oddomorphism $f:G\xrightarrow{\text{odd}}K_{n}$, two vertices $u_{1},u_{2}$ such that $f(u_{1})=f(u_{2})$ and a (possibly empty) set $\mathcal{P}$ of pairwise edge-disjoint $(u_{1},u_{2})$-paths, such for each path $\in\mathcal{P}, PP$ is $2$-coloured under $f$, we define the $\mathcal{P}$-merger of $u_{1}$ and $u_{2}$ to be the graph $G^{\prime}$ obtained by identifying $u_{1}$ and $u_{2}$ into a single vertex $u^{\prime}$ whose neighbourhood is the symmetric difference of the neighbourhoods of $u_{1}$ and $u_{2}$, and then deleting all of the edges from the paths in $\mathcal{P}$.
$u_{1}$$u_{2}$$\mathcal{P}$$N(u_{1})-N(u_{2})$$N(u_{1})\cap N(u_{2})$$N(u_{2})-N(u_{1})$$\mathcal{P}$$u^{\prime}$$N(u_{1})-N(u_{2})$$N(u_{1})\cap N(u_{2})$$N(u_{2})-N(u_{1})$
Figure 1: An example of a $\mathcal{P}$-merger of $u_{1}$ and $u_{2}$. Note that the paths in $\mathcal{P}$ may also be in the union of $N(u_{1})$ and $N(u_{2})$.
In the event $\mathcal{P}$ is empty, we simply call this the merger of $u_{1}$ and $u_{2}$. See Figure 1 for an illustration.
Lemma 2.4.
Suppose that $f:G\xrightarrow{\text{odd}}K_{n}$ and $G^{\prime}$ is obtained from $G$ by the $\mathcal{P}$-merger of vertices $u_{1}$ and $u_{2}$ into $u^{\prime}.$ Let $f^{\prime}:V(G)\to V(K_{n})$ be the map where $f^{\prime}(u^{\prime})=f(u_{1})=f(u_{2})$ and, for all $v\in V(G)-\{u_{1},u_{2}\}$, we have $f^{\prime}(v)=f(v)$. Then we have $f^{\prime}:G\xrightarrow{\text{odd}}K_{n}$.
Proof.
Assume without loss of generality that $V(K_{n})=\{1,\dots,n\}$ and for $c\in V(K_{n})$, let $C_{c}=f^{-1}(\{c\})$ and $C_{c}^{\prime}=(f^{\prime})^{-1}(\{c\})$. Observe that $f^{\prime}$ is a homomorphism as $u_{1}$ and $u_{2}$ have the same colour under $f$. Let $c=f\!\left(u^{\prime}\right)$.
First consider a vertex $v\in V\!\left(G^{\prime}\right)-u^{\prime}$ and let $P_{1},\dots,P_{\ell}$ denote the paths in $\mathcal{P}$ containing $v$ (where possibly $\ell=0$). Let $c^{\prime}=f(v)$ (where possibly $c^{\prime}=c$). If $c^{\prime}\neq c$, then as each path in $\mathcal{P}$ is $2$-coloured and has endpoints $u_{1}$ and $u_{2}$, every neighbour of $v$ in $P_{1},\ldots,P_{\ell}$ is coloured $c$. Thus $|N_{G^{\prime}}(v)\cap C_{c}^{\prime}|=|N_{G}(v)\cap C_{c}|-(2\ell-2b)$, where $b\in\{0,1\}$ and $b=1$ if and only if $v$ is adjacent to both $u_{1}$ and $u_{2}$ and the path $u_{1},v,u_{2}$ is not in $\mathcal{P}$. Therefore $|N_{G^{\prime}}(v)\cap C_{c}^{\prime}|$ has the same parity as $|N_{G}(v)\cap C_{c}|$.
Now assume that $c^{\prime}=c$. Then $v$ is adjacent to neither $u_{1}$ nor $u_{2}$. For each colour $c^{\prime\prime}$, let $q_{c^{\prime\prime}}$ be the number of paths in $\mathcal{P}$ that contain $v$ and a vertex coloured $c^{\prime\prime}$. Then $|N_{G^{\prime}}(v)\cap C_{c^{\prime\prime}}^{\prime}|=|N_{G}(v)\cap C_{c^{\prime\prime}}|-2q_{c^{\prime\prime}}$, hence $|N_{G^{\prime}}(v)\cap C_{c}^{\prime}|$ has the same parity as $|N_{G}(v)\cap C_{c}|$.
Now consider $u^{\prime}$. Fix a colour $c^{\prime}\neq c$. Let $q_{c^{\prime}}$ be the number of paths in $\mathcal{P}$ which contain a vertex coloured $c^{\prime}$, and let $p_{c^{\prime}}$ be the number of vertices in the common neighbourhood of $u_{1}$ and $u_{2}$ coloured $c^{\prime}$ which do not appear in a path in $\mathcal{P}$. Then
$|N_{G^{\prime}}(u^{\prime})\cap C_{c^{\prime}}^{\prime}|=|N_{G}(u_{1})\cap C_{c^{\prime}}|+|N_{G}(u_{2})\cap C_{c^{\prime}}|-2q_{c^{\prime}}-2p_{c^{\prime}}$
Therefore if $u_{1}$ and $u_{2}$ are both $f$-even or both $f$-odd, then $u^{\prime}$ is $f^{\prime}$-even, and if exactly one of $u_{1}$ and $u_{2}$ is $f$-even, then $u^{\prime}$ is $f^{\prime}$-odd.
Therefore all vertices in $G^{\prime}$ are either $f^{\prime}$-even or $f^{\prime}$-odd. The last remaining condition is that there are an odd number of $f^{\prime}$-odd vertices of each colour class. Other than colour $c$, the number of $f$-odd vertices remains the same in $f$ and $f^{\prime}$. For colour $c$, if both $u_{1}$ and $u_{2}$ were $f$-odd, then $u^{\prime}$ is $f$-even, and there was at least three and an odd number of $f$-odd vertices of colour $c$ in $f$. Therefore there is an odd number of $f$-odd vertices of colour $c$ under $f^{\prime}$, and hence $f^{\prime}$ is an oddomorphism. ∎
We also observe that removing bi-coloured cycles preserves oddomorphisms. For the next lemma, we use the notation that given two disjoint sets of vertices, $A,B\subseteq V(G)$ the graph $G[A,B]$ is the graph containing all edges with one endpoint in $A$ and one endpoint in $B$.
Lemma 2.5.
Let $f:G\xrightarrow{\text{odd}}K_{n}$ and let $C_{i}$ and $C_{j}$ be two colour classes, $i\neq j$. Let $C$ be a cycle in $G[C_{i},C_{j}]$. Then $f:G-E(C)\xrightarrow{\text{odd}}K_{n}$.
Proof.
As we only deleted edges, $f$ is a homomorphism from GE(C)G-E(C) to $K_{n}$. The parity of every vertex remains the same in GE(C)G-E(C) as $G$, since $C$ is $2$-coloured, so the number of neighbours coloured $i$ or $j$ changes either by 0 or $2$. It follows that $f:G-E(C)\xrightarrow{\text{odd}}K_{n}$. ∎
In particular, Lemma 2.5 implies that if an oddomorphism is edge-minimal, it is a so-called acyclic colouring.
3 Large oddomorphism implies large complete immersion
For ease, given an $(x,y)$-path $P$, we say that to split a path $P$ is to iteratively split off edges of $P$ to create the edge $xy$.
Suppose, for a contradiction, that Theorem 1.8 is false, and consider a vertex-minimal, and subject to that, edge-minimal counterexample. By Lemma 2.5, we may assume that $G[C_{i},C_{j}]$ is a forest for all $i\neq j$.
Claim.
Let $C_{i}$ and $C_{j}$ be two distinct colour classes of $f$. Then there is no path of length at least $2$ from an odd vertex in $C_{i}$ to an odd vertex in $C_{j}$ in $G[C_{i},C_{j}]$.
[…]
Thus we have that for every two colour classes $C_{i}$ and $C_{j}$, the graph $G[C_{i},C_{j}]$ is a forest. Note that a vertex in $C_{i}\cup C_{j}$ has odd degree in $G[C_{i},C_{j}]$ if and only if it is $f$-odd. It follows that we can decompose $G[C_{i},C_{j}]$ into (edge-disjoint) paths such that each path has two $f$-odd vertices in $G[C_{i},C_{j}]$ as its endpoints.
Let $G^{\prime}$ be the graph obtained by splitting off each such path for all pairs $i,j$ with $i\neq j$ and removing all even degree vertices. Let $G^{\prime\prime}$ be the simplification of $G^{\prime}$ – that is, $G^{\prime\prime}$ is the graph obtained from $G^{\prime}$ by removing all parallel edges. If $G^{\prime\prime}$ has minimum degree at least $7t+7$, then by Theorem 2.1, $G^{\prime\prime}$ has a $K_{t}$-immersion, and as $G$ has $G^{\prime\prime}$ as an immersion, it follows that $G$ has a $K_{t}$-immersion. Thus we can assume that $G^{\prime\prime}$ has minimum degree less than $7t+7$. Now by Proposition 2.2, each odd vertex has at least one neighbour in each colour other than itself. Thus each odd vertex in $G^{\prime}$ has degree at least ${t\choose 2}(7t+7)-1$ in $G^{\prime}$. As $G^{\prime\prime}$ has minimum degree less than $7t+7$, this implies that there are two vertices $x,y$ such that there are at least $t\choose 2$} parallel edges between them, as otherwise vertices could have degree at most $({t\choose 2}-1)(7t+7)$.
Let $x,y$ be two such vertices. By construction, there are at least $t\choose 2$} edge-disjoint paths $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{{t\choose 2}}\}$ from $x$ to $y$ such that each path is $2$-coloured under $f$.
By the claim if two odd vertices have different colours, either they are adjacent or there is no path between them using just two colours. By this observation and our construction it follows that $f(x)=f(y)$.
Let $G^{*}$ denote the $\mathcal{P}$-merger of $x$ and $y$ and let $f^{*}$ be the $\binom{t}{2}(7t+7)$-oddomorphism of $G^{*}$ as in Lemma 2.4.
By the minimality of $G$, $G^{*}$ has a $K_{t}$-immersion. We claim that this implies that $G$ has a $K_{t}$-immersion. To see this, we use the definition of immersions given in Section 2.3. Let $t_{1},\ldots,t_{t}$ be the branch vertices of the $K_{t}$-immersion in $G^{*}$ and, for each $1\leq i<j\leq t$, let $Q_{ij}$ be the path from $t_{i}$ to $t_{j}$ (such that $Q_{ij}$’s are mutually edge-disjoint). Observe that if $Q_{ij}$ uses the merged vertex more than once, $Q_{ij}$ in fact contains a cycle. Thus each path $Q_{ij}$ uses the merged vertex at most once. Therefore we can extend this immersion to an immersion of $G$ in the following manner. For each $Q_{ij}$ that uses the merged vertex, we use one of the paths $P_{1},\ldots,P_{{t\choose 2}}$ to extend $Q_{i,j}$ to a path in $G$. As there are at most $t\choose 2$} paths $Q_{ij}$, we can do this such that each $P_{i}$ is used at most once, and thus we have a $K_{t}$-immersion in $G$.
4 A tight bound for oddomorphisms and treewidth
To prove Theorem 1.10, suppose towards a contradiction that $\operatorname{tw}(G)\leq t$, $G$ has a $(t+2)$-oddomorphism $f$, and subject to these conditions, $G$ minimizes $|V(G)|+|E(G)|$. As removing edges does not increase the treewidth, Lemma 2.5 implies that for any two colour classes $C_{i},C_{j}$, we have that $G[C_{i},C_{j}]$ is acyclic. Let $(T,\{\beta_{t}\}_{t\in V(T)})$ be a tree decomposition of $G$ of width at most $t$. Fix a vertex $r\in V(T)$ and view $T$ as a tree rooted at $r$.
For every $f$-odd vertex $u$ in $G$, let $D(u,r)$ be the shortest distance in $T$ from $r$ to a bag containing $u$. Let $w$ be an $f$-odd vertex which maximizes $D(u,r)$ over all $f$-odd vertices $u$ (picking arbitrarily if there are more than one). Let $\beta_{v}$ be the bag which attains $D(w,r)$. Note that $v\neq r$ because there are at least $t+2$ ff-odd vertices and $|\beta_{r}|\leq t+1$.
As $w$ is $f$-odd, for each of the $t+1$ colours distinct from $f(w)$, $w$ has a neighbour of that colour. Since $w\in\beta_{v}$ and $|\beta_{v}|\leq t+1$, there exists a colour $c^{\prime}$ such that no vertex in $\beta_{v}$ has colour $c^{\prime}$. Let $w^{\prime}$ be a neighbour of $w$ of colour $c^{\prime}$. Note that if $w^{\prime}$ appears in a bag of the component of TvT-v containing $r$, then $w$ also appears in that bag by the definition of tree decompositions, contradicting our choice of $\beta_{v}$. Hence, $w^{\prime}$ appears in a bag of a component of TvT-v not containing $r$. Let $c=f(w)$.
Observe that in $G[C_{c},C_{c^{\prime}}]$ there is a path starting at $w$, going through $w^{\prime}$ and ending in some $f$-odd vertex $w^{\prime\prime}$. Let $P$ be such a path. If PwP-w intersects $\beta_{v}$, then by our choice of the colour $c^{\prime}$, there is a vertex $q\in\beta_{v}$ such that $q\neq w$ and $q$ has colour $c$. In this case, consider the merger of $q$ and $w$ and let $G^{\prime}$ be the resulting graph. Then $|V(G^{\prime})|<|V(G)|$ and, by Lemma 2.4, $G^{\prime}$ admits a $(t+2)$-oddomorphism. For each $t\in V(T)$, let $\beta_{t}^{\prime}$ be the bag obtained from $\beta_{t}$ by replacing $q$ and $w$ (if they appear) with the merged vertex. Since $q$ and $w$ are in a common bag of $(T,\{\beta_{t}\}_{t\in V(T)})$, we have that $(T,\{\beta_{t}^{\prime}\}_{t\in V(T)})$ is a tree decomposition of $G^{\prime}$ of width $t$, contradicting the minimality of the counterexample $G$.
Thus $\beta_{v}$ has no vertex of $P$ except $w$. Hence the $f$-odd vertex $w^{\prime\prime}$ appears only in bags of components of TvT-v not containing $r$, contradicting our choice of $w$.
5 Homomorphism counting and oddomorphisms
First we claim that to prove Theorem 1.3 it suffices to prove the following:
Theorem 5.1.
For any $n\in\mathbb{N}$, there exists non-isomorphic graphs $H$ and $H^{\prime}$ such that $\hom(F,H)\neq\hom(F,H^{\prime})$ implies that $F$ contains a $K_{n}$-immersion.
The next proof is very similar to a theorem in [22].
Theorem 5.2.
Theorem 5.1 and Theorem 1.3 are equivalent.
Proof.
First suppose Theorem 5.1 is true and we are to prove Theorem 1.3. Let $\mathcal{F}$ be a proper immersion-closed and union-closed class. As $\mathcal{F}$ is proper and union closed, there is some $n$ such that $K_{n}$ is not in $\mathcal{F}$. By Theorem 5.1, there are non-isomorphic graphs $H$ and $H^{\prime}$ such that $\hom(F,H)\neq\hom(F,H^{\prime})$ implies that $F$ contains $K_{n}$ as an immersion. But then for all $K\in\mathcal{F}$, we have $\hom(K,H)=\hom(K,H^{\prime})$, as $K_{n}\not\in\mathcal{F}$. Thus $H=_{\mathcal{F}}H^{\prime}$, implying that $_{\mathcal{F}}$ is not isomorphism.
Now suppose that Theorem 1.3 is true. Let $n\in\mathbb{N}$ be given, and let $\mathcal{F}$ be the class of graphs with no $K_{n}$-immersion. By Theorem 1.3, there exist non-isomorphic graphs $H,H^{\prime}$ such that for all $K\in\mathcal{F}$, we have $\hom(K,H)=\hom(K,H^{\prime})$. Hence, if $\hom(F,H)\neq\hom(F,H^{\prime})$, then $F\not\in\mathcal{F}$, which implies that $F$ has a $K_{n}$-immersion by definition of $\mathcal{F}$. ∎
Now our goal is to prove Theorem 5.1.
We need to utilize a theorem from [22], and as such we need to extend the notion of oddomorphism from cliques to all graphs.
Definition 5.3.
Let $f:V(G)\rightarrow H$ be a homomorphism. A vertex $u\in V(G)$ is $f$-odd (respectively, $f$-even) if $N_{G}(v)$ contains an odd (respectively, even) number of vertices of each colour in $S\setminus\{f(v)\}$for every $v\in V(H)\cap N_{H}(f(u))$, the number of vertices $w$ in $N_{G}(u)$ such that $f(w)=v$ is odd. Similarly, $u$ is $f$-even if for every $v\in V(H)\cap N(f(u))$, the number of vertices $w$ in $N_{G}(u)$ such that $f(w)=v$ is even.
Definition 5.4.
Given graphs $G$ and $H$ and a homomorphism $f:V(G)\to V(H)$, we say that $f$ is an oddomorphism from $G$ to $H$, denoted $G\xrightarrow{\text{odd}}H$, if:
  1. 1.
    Every vertex of $G$ is either $f$-even or $f$-odd.
  2. 2.
    For every vertex $v\in V(H)$, there is an odd number of vertices in $G$ whose image under $f$ is $v$ and are $f$-odd.
Definition 5.5.
Let $f:G\to F$. We say $f$ is a weak oddomorphism if there is a (not necessarily induced) subgraph $G^{\prime}$ of $F$ such that the restriction of $f$ to $G^{\prime}$ is an oddomorphism from $G^{\prime}$ to $F$.
Theorem 5.6 (Theorem 3.13, [22]).
Let $G$ be a connected graph. Given $G$, there are non-isomorphic graphs $G_{0}$ and $G_{1}$ constructed from $G$ such that $\hom(F,G_{0})\geq\hom(F,G_{1})$ with strict inequality if and only if there exists a weak oddomorphism from $F$ to $G$. Moreover, if such a weak oddomorphism $\phi$ does exist, then there is a connected subgraph $F^{\prime}$ of $F$ such that $\phi|_{V(F^{\prime})}$ is an oddomorphism from $F^{\prime}$ to $G$.
We note that in Theorem 3.13 in [22], the fact that $G_{0}$ and $G_{1}$ are non-isomorphic is not stated in the theorem. However this follows from Lemma 3.14 in [22] and Theorem 1.1.
Proof of Theorem 5.1.
Let $n\in\mathbb{N}$ be given. Let $G=K_{{n\choose 2}(7n+7)}$. By Theorem 5.6, there are non-isomorphic graphs $G_{0},G_{1}$ such that $\hom(F,G_{0})\geq\hom(F,G_{1})$ and strict inequality holds if and only if $F$ has a weak oddomorphism to $G$. As $G_{1}$ and $G_{0}$ are not isomorphic, by Theorem 1.1 there is a graph $F$ such that $\hom(F,G_{0})\neq\hom(F,G_{1})$. For all such graphs $F$, we have that $F$ admits a weak oddomorphism to $G$. Thus there is a subgraph $F^{\prime}$ of $F$ such that $F^{\prime}\xrightarrow{\text{odd}}G$. But by Theorem 1.8, this implies $F^{\prime}$ has a $K_{t}$-immersion, and hence $F$ has a $K_{t}$-immersion, thus implying Theorem 5.1. ∎
We note that we could perform a similar proof to show that if $\mathcal{M}$ is the class of graphs with treewidth at most $k$, then $_{\mathcal{M}}$ is not isomorphism, however this is already known [20].
Acknowledgments.
Benjamin Moore thanks Sabrina Lato and Leqi (Jimmy) Zhu for inspiring discussions on oddomorphisms. Youngho Yoo and Benjamin Moore thank the Pacific Institute of Mathematics for travel support to the University of Manitoba. Benjamin Moore thanks the Chilean research agency ANID for travel support.